OpenStudy (anonymous):

What is the equation of a line passing through the point (6,1) and a parallel to a line whose eqution is 3x=2y+4?

OpenStudy (anonymous):

y=3/2x-2 so slope is 3/2 y-1=3/2(x-6) y=3/2x-8

OpenStudy (anonymous):

y = m x + b Solve 3x = 2y +4 for y. The result should have the form shown above and should provide the value of m. The new line passes through the point (6,1) and therefore provides values for the new x and y. Plug the value of m from paragraph 1 into the expression below and solve the result for b. 1 = m(6) + b